OpenStudy (anonymous):

Find the width of a rectangle if the length is 7cm and perimeter is 35cm. (1 points) 10cm 10.5cm 7.5cm 5cm

Mehek (mehek14):

ok there are 2 lengths and 2 widths you know the length is 7 what is 7*2

OpenStudy (anonymous):

P = 2l + 2w 35 = 2 times 7 + 2w

OpenStudy (anonymous):

35 = 14 + 2w

Mehek (mehek14):

@WhiteFangIII

OpenStudy (anonymous):

14

Mehek (mehek14):

now what is 35 - 14

OpenStudy (anonymous):

35 - 14 = 2w

OpenStudy (anonymous):

21 = 2w

OpenStudy (anonymous):

w = 21/2

OpenStudy (anonymous):

w = 10.5cm

OpenStudy (anonymous):

35-14=21

Mehek (mehek14):

ok that is the value of 2 widths you only need one so what is 21/2

OpenStudy (anonymous):

10.5

OpenStudy (anonymous):

CORRECT!

Mehek (mehek14):

that is your answer